OpenStudy (jazzyfa30):

solve each equation. give an exact solution and a four-decimal place approximation. 7. 5^(2x)=12 8, ln(x+3)=2 9. 4^(x-2)=3 10. 2^(x-3=6^(1-2x)

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

if u see exponents in the equation, and if u dont see logs anywhere, then taking log both sides is a good idea

ganeshie8 (ganeshie8):

\(\large 5^{2x}=12\) take log both sides \(\large \log( 5^{2x})= \log (12)\)

ganeshie8 (ganeshie8):

next, u need to knw few log properties : \(\large \log_a (b)^m = m ~\log_a b\)

ganeshie8 (ganeshie8):

use it

OpenStudy (jazzyfa30):

how do i break down 12

ganeshie8 (ganeshie8):

\(\large 5^{2x}=12\) take log both sides \(\large \log( 5^{2x})= \log (12)\) \(\large 2x~\log( 5)= \log (12)\)

ganeshie8 (ganeshie8):

divide both sides by \(\log (5)\)

ganeshie8 (ganeshie8):

\(\large 5^{2x}=12\) take log both sides \(\large \log( 5^{2x})= \log (12)\) \(\large 2x~\log( 5)= \log (12)\) \(\large 2x = \frac{\log (12)}{\log(5)}\)

ganeshie8 (ganeshie8):

use ur calculator and find out both logs on right hand side

ganeshie8 (ganeshie8):

and divide

OpenStudy (jazzyfa30):

log12 is 1.079181246 and log5 is .6989700043

ganeshie8 (ganeshie8):

yes, divide them

OpenStudy (jazzyfa30):

log 12/log5 is 1.543959311

ganeshie8 (ganeshie8):

yes ! so, \(\large 5^{2x}=12\) take log both sides \(\large \log( 5^{2x})= \log (12)\) \(\large 2x~\log( 5)= \log (12)\) \(\large 2x = \frac{\log (12)}{\log(5)}\) \(\large 2x = 1.543959311\)

ganeshie8 (ganeshie8):

solve \(x\)

OpenStudy (jazzyfa30):

1.543959311/2 is .7719796555

ganeshie8 (ganeshie8):

correct ! round it as asked...

OpenStudy (jazzyfa30):

1

ganeshie8 (ganeshie8):

whut

ganeshie8 (ganeshie8):

.7719796555 rounded to four decimal places is .7720

OpenStudy (jazzyfa30):

ohhh ok i didnt learn that

OpenStudy (jazzyfa30):

its time for lunch when i come back i will let you know so we can finish and thank you so much

ganeshie8 (ganeshie8):

enjoy ur dinner !

ganeshie8 (ganeshie8):

lol here its dinner time in india :)

OpenStudy (jazzyfa30):

im back

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

hey sorry, missed this...

ganeshie8 (ganeshie8):

for next question, base of ln is always \(e\)

ganeshie8 (ganeshie8):

\(\large \ln(x+3)=2\) is same as \(\large \ln_e(x+3)=2\)

ganeshie8 (ganeshie8):

change it to exponent form...

OpenStudy (jazzyfa30):

e^2=x+3

ganeshie8 (ganeshie8):

yes, solve \(x\)

ganeshie8 (ganeshie8):

and keep in mind, "e" is just a NUMBER, like "pi"

OpenStudy (jazzyfa30):

x=1

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

e^2 = x + 3 subtract 3 both sides e^2 - 3 = x

OpenStudy (jazzyfa30):

5.389

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=e%5E2-3

ganeshie8 (ganeshie8):

you should get x = 4.3891

OpenStudy (jazzyfa30):

i see it now